# Lower Bound for The Number of Zeros in The Character Table of The Symmetric Group

### v3

@@ Abstract.
$Z(N)\geq\frac{2\,p(N)^{2}}{\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right),$
@@ 1. Introduction
For any two partitions $\lambda$ and $\mu$ of a positive integer $N$, let $\chi_{\lambda}(\mu)$ denote the value of the irreducible character of the symmetric group $S_{N}$ associated with $\lambda$, evaluated in the conjugacy class of elements whose cycle type is determined by $\mu$. By the Murnaghan-Nakayama rule [4], it is known that irreducible characters are integer-valued functions, and the number of irreducible characters of $S_{N}$ is equal to $p(N)$, the number of partitions of $N$. In this article, we study the zeros of the character values. Although linear characters never take the value zero, Burnside’s classical result [1] establishes that every non-linear irreducible character must vanish at some group element. Miller [7] proved that if one chooses an irreducible character of $S_{N}$ uniformly at random and selects a random element from $S_{N}$ uniformly, then the probability that the character value is zero approaches $1$ as $N\to\infty$. However, this result does not estimate the number of zeros in the character table of $S_{N}$ since the character values are distributed over the conjugacy classes, rather than individual elements of $S_{N}$. Let $Z(N)$ be the number of zeros in the character table of the symmetric group $S_{N}$. Miller [7, 8] introduced the problem of determining the asymptotic behavior of $Z(N)$. Due to the rapid growth of $p(N)$, computation of $Z(N)$ is challenging. Recently, Miller and Scheinerman [9] conducted a large-scale Monte Carlo simulation to determine the density of zeros in the character table of $S_{N}$ for large values of $N$, leading to the following conjecture:
@@
Peluse [11] proved that the proportion of zeros in the character table of $S_{N}$ is at least $M/\log N$ for some positive constant $M$. Here, we aim to determine an explicit value for $M$. To achieve this, we need a lower bound for the number of $t$-core partitions $c_{t}(N)$. In a recent paper [10], Morotti proved such a lower bound, which Peluse and Soundararajan utilized in [12]. Using Morotti’s bound, one can deduce the following inequality:
$Z(N)\geq\frac{p(N)^{2}}{2\log N}\left(1+O\left(\frac{1}{\log N}\right)\right).$
In [13], Peluse and Soundararajan mention that $Z(N)\geq\frac{2p(N)^{2}}{\log N}$, without a proof. In this article, we prove the above lower bound.
@@
$Z(N)\geq\frac{2p(N)^{2}}{\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right).$
As the above bound matches with the Miller and Scheinerman conjecture, it may be difficult to improve this bound further.
@@
where $p_{t}(N)$ denotes the number of partitions of $N$ into parts of size at most $t$. Here, we attempt to obtain an exact order of the above sum. However, this requires asymptotic estimates for $c_{t}(N)$ and $p_{t}(N-t)$. We take the asymptotic formula for $c_{t}(N)$ from a recent paper of Tyler [15] and the asymptotic formula for $p_{t}(N-t)$ from Erdös and Lehner [3]. We will see later in the proof that Tyler’s formula plays an important role as it gives asymptotic bound for $c_{t}(N)$ as $t$ and $N$ both varies. In the later part of the proof, we treat the above sum in different ranges of $t$ to obtain an optimal bound.
@@
McSpirit and Ono [6] proved the following result for primes $t\geq 5$:
@@
Let $N$ be a large positive integer, and $t\leq N$. Then we have the following results:
(i) For $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$ and for any $0<\epsilon<1$,
@@
(ii) For $\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}<t<\frac{2\sqrt{6N}}{\sqrt{6/\pi-1}},$
@@
$Q_{t}(N)=\frac{2\sqrt{\pi}\exp\left(\frac{t}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}.$
(iii) For $\frac{2\sqrt{6N}}{\sqrt{6/\pi-1}}\leq t$,
$Z_{t}(N)\geq\frac{p(N)^{2}}{\exp\left(1.00873t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)+\frac{2\pi}{\sqrt{6}}\frac{t}{\sqrt{N-t}+\sqrt{N}}\right)}\left(1+O\left(\frac{t}{N}\right)\right).$
@@ 2. Acknowledgments
We extend our heartfelt thanks to Prof. Ben Kane for his insightful and valuable suggestions. J. Barman is deeply thankful to the University Grants Commission (UGC), India, for their invaluable support through the Fellowship Programme. K. Mahatab is supported by the DST INSPIRE Faculty Award Program and grant no. DST/INSPIRE/04/2019/002586.
@@ 3. Preliminaries
If we consider the parttions $\mu$ having largest part $t$ and the $t$-core partitions $\lambda$, then by the above lemma $\chi_{\lambda}(\mu)=0.$ This set of pairs $(\lambda,\mu)$ gives $c_{t}(N)p_{t}(N-t)$ zeros. We also notice that the set of partitions $\mu$ of $N$ having largest part $t_{1}$ is disjoint from the set of partitions of $N$ having largest part $t_{2}$. So the number of zeros in the character table is at least
@@
Next, we simplify the Tyler’s formula for $c_{t}(N)$, which requirs the following notations.
@@
where $z=x+iy$. In [15], Tyler defines the following functions:
$\mu_{k}(z)=-\frac{z^{k+1}}{2\pi i}\left(\frac{d}{dz}\right)^{k}\log\eta(z),$
$f_{t}(z)=\frac{\eta(tz)^{t}}{\eta(z)}.$
@@
For $x\in\mathbb{R}$ and $y\geq\frac{\sqrt{3}}{2}$,
$\eta(iy)=\exp\left(-\frac{\pi y}{12}-ve^{-2\pi y}\right)$
with $1<v<1.00873.$
@@
Let $x\in\mathbb{R}$ and $y$ be a small positive real number. Then
$\eta(iy)=y^{-\frac{1}{2}}\exp\left(-\frac{\pi}{12y}-ve^{-\frac{2\pi}{y}}\right)$
with $1<v<1.00873$.
@@
$\eta\left(-\frac{1}{z}\right)=\sqrt{-iz}\,\eta(z).$
@@
Using the above result in Lemma 3.3, we conclude the proof. ∎
Lemma 3.5 ([15, Lemma 3.2]).
Let $\mu_{2}$ and $y$ be defined as before. Then for any positive integer $t$, the following inequalities hold:
$(i)$ If $ty<1$, then
$\frac{4\pi}{ty-y}<\frac{1}{\mu_{2}(iy)-\mu_{2}(ity)}<\frac{8\pi}{ty-y}.$
$(ii)$ If $ty\geq 1$, then
@@
Below, we simplify Tyler’s bound [15] for $c_{t}(N)$ in different ranges of $t$.
@@
Let $N$ be a large positive integer and $t\leq N$.
(i) For $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$ and for any $0<\epsilon<1$, we have
@@
(ii) For $\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}<t<\frac{2\sqrt{6N}}{\sqrt{6/\pi-1}},$ we have
$c_{t}(N)\geq\frac{2\sqrt{\pi}\exp\left(\frac{t}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}\left(1+O\left(t^{-1}\right)\right).$
(iii) For $\frac{2\sqrt{6N}}{\sqrt{6/\pi-1}}\leq t$, we have
$c_{t}(N)\geq p(N)\exp\left(-1.00873t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
@@
$(i)$ From Theorem 1.4 of [15], it follows that for $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$ and $0<\epsilon<1$, the following holds:
@@
Substituting this into (3.1), we obtain
@@
$(ii)$ From Theorem 1.4 of [15], we know that
(3.2) $c_{t}(N)=\frac{y^{\frac{3}{2}}\exp\left(2\pi y\left(N+\frac{t^{2}-1}{24}\right)\right)f_{t}(iy)}{\sqrt{\mu_{2}(iy)-\mu_{2}(ity)}}\left(1+O\left(t^{-1}\right)\right),$
and also, $y$ satisfies the equation
(3.3) $\frac{\mu_{1}(ity)-\mu_{1}(iy)}{y^{2}}=N+\frac{t^{2}-1}{24},$
and $y$ lies in the range
$\frac{t-1}{4\pi\left(N+\frac{t^{2}-1}{24}\right)}<y<\frac{1}{\frac{3}{\pi}+\sqrt{24N-1+\frac{9}{\pi^{2}}}}.$
For the given range of $t$, we observe that $ty\leq 1$. Using Lemma 3.4 and Lemma 3.5 in (3.2), we obtain
@@
We may verify that
$y=\frac{t}{4\pi\left(N+\frac{t^{2}-1}{24}\right)}$
is a feasible solution to (3.3). We obtain
$c_{t}(N)\geq\frac{2\sqrt{\pi}\exp\left(\frac{t}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}\left(1+O\left(t^{-1}\right)\right).$
$(iii)$ From Theorem 1.4 of [15], we have the following identity:
(3.4) $c_{t}(N)=\frac{y^{\frac{3}{2}}\exp\left(2\pi y\left(N+\frac{t^{2}-1}{24}\right)\right)f_{t}(iy)}{\sqrt{\mu_{2}(iy)-\mu_{2}(ity)}}\left(1+O\left(N^{-\frac{1}{2}}\right)\right),$
Note that for $ty\geq 1$, $y=\frac{1}{\sqrt{24N}}$ is a feasible solution to (3.3). Now, employing $y=\frac{1}{\sqrt{24N}}$ in the range $\frac{2\sqrt{6N}}{\sqrt{6/\pi-1}}\leq t$, we get $ty\geq\frac{\sqrt{3}}{2}$. Then, using the Lemmas 3.3, 3.4 and 3.5 in the formula (3.4), we obtain
$c_{t}(N)$ $\geq\sqrt{12}y^{2}\exp\left(y\left(2\pi N-\frac{\pi}{12}\right)+\frac{\pi}{12y}-1.00873t\exp(-2\pi yt)+e^{-\frac{2\pi}{y}}\right)$
$\hskip 233.3125pt\times\left(1+O\left(N^{-\frac{1}{2}}\right)\right)$
$=\frac{\exp\left(\frac{2\pi}{\sqrt{6}}\sqrt{N}\right)}{4\sqrt{3}N}\exp\left(-1.00873t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
Therefore, we conclude that
$c_{t}(N)\geq p(N)\exp\left(-1.00873t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
The following corollary is required in the proof of Theorem 1.2.
Corollary 3.7.
Let $\frac{\sqrt{6}}{2\pi}\sqrt{N}\log N<t\leq N$. Then
$c_{t}(N)\geq p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\log N\right)\right).$
Proof.
From the proof of Lemma 3.3, we can write
$\eta(ity)=\exp\left(-\frac{\pi ty}{12}-ve^{-2\pi ty}\right),$
where $1<v<\alpha=\frac{1}{(1-\exp(-2\pi yt))^{2}}$. One can easily check that for the above range of $t$, $\alpha=1+O\left(N^{-\frac{1}{2}}\right)$. From the proof of $(iii)$ of Proposition 3.6, we can write $c_{t}(N)$ as
@@
$\hskip 233.3125pt\times\left(1+O\left(N^{-\frac{1}{2}}\right)\right),$
which simplifies to
$c_{t}(N)\geq p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\log N\right)\right)$
in the given range of $t$. ∎
In 1941, Erdös and Lehner [3] proved the following result for $p_{t}(N)$ without an error term. We give a sketch of the proof of this result, including an explicit error term.
@@
where $C=2\pi/\sqrt{6}$ and $x\ll N^{\frac{1}{4}-\epsilon},\ \epsilon>0$.

### v4

@@ Abstract.
$Z(N)\geq\frac{2\,p(N)^{2}}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right),$
@@
Key words and phrases:
Zeros in Character Table, Symmetric Groups, Partition Function, t-core Partitions
@@ 1. Introduction
For any two partitions $\lambda$ and $\mu$ of a positive integer $N$, let $\chi_{\lambda}(\mu)$ denote the value of the irreducible character of the symmetric group $S_{N}$ associated with $\lambda$, evaluated in the conjugacy class of elements whose cycle type is determined by $\mu$. By the Murnaghan-Nakayama rule [4], it is known that irreducible characters are integer-valued functions, and the number of irreducible characters of $S_{N}$ is equal to $p(N)$, the number of partitions of $N$. In 1918, Hardy and Ramanujan [5] used the circle method to prove the celebrated asymptotic formula for $p(N)$:
$p(N)\sim\frac{1}{4N\sqrt{3}}\exp\bigg(\frac{2\pi}{\sqrt{6}}\sqrt{N}\bigg).$
A more usable and explicit version of the above result, proved by Rademacher [13], gives:
(1.1) $p(N)=\frac{1}{4N\sqrt{3}}\exp\bigg(\frac{2\pi}{\sqrt{6}}\sqrt{N}\bigg)\bigg(1+O(N^{-\frac{1}{2}})\bigg),$
which we will use frequently. In this article, we study the zeros of the character values. Although linear characters never take the value zero, Burnside’s classical result [1] establishes that every non-linear irreducible character must vanish at some group element. A statistical estimate proved by Miller [9] states that if one chooses an irreducible character of $S_{N}$ uniformly at random and selects a random element from $S_{N}$ uniformly, then the probability that the character value is zero approaches $1$ as $N\to\infty$. However, this result does not estimate the number of zeros in the character table of $S_{N}$ since the character values are distributed over the conjugacy classes, rather than individual elements of $S_{N}$. Note that the character table of $S_{N}$ has $p(N)^{2}$ entries. Let $Z(N)$ be the number of zeros in the character table of the symmetric group $S_{N}$. Miller [9, 10] introduced the problem of determining the asymptotic behavior of $Z(N)$. Due to the rapid growth of $p(N)$, computation of $Z(N)$ is challenging. To tackle this problem, Miller and Scheinerman [8] conducted a large-scale Monte Carlo simulation to determine the density of zeros for large values of $N$, leading to the following conjecture:
@@
Peluse [12] proved that the proportion of zeros in the character table of $S_{N}$ is at least $A/\log N$ for some positive constant $A$. Later, in [11], Peluse and Soundararajan remarked that $Z(N)\geq\frac{2p(N)^{2}}{\log N}$, without any details. In this article, we prove the above lower bound.
@@
$Z(N)\geq\frac{2p(N)^{2}}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
If Conjecture 1.1 is true, then the main term of our lower bound is optimal.
@@
where $p_{t}(N)$ denotes the number of partitions of $N$ into parts of size at most $t$. Here, we attempt to obtain an exact order of the above sum. However, this requires asymptotic estimates for $c_{t}(N)$ and $p_{t}(N-t)$. We take the asymptotic formula for $c_{t}(N)$ from the recent work of Tyler [14] and the asymptotic formula for $p_{t}(N-t)$ from Erdös and Lehner [3]. We will see later in the proof that Tyler’s formula plays an important role as it gives an asymptotic bound for $c_{t}(N)$ when $t$ and $N$ both vary. In the later part of the proof, we treat the above sum in different ranges of $t$ to obtain an optimal bound.
@@
McSpirit and Ono [7] proved the following result for fixed primes $t\geq 5$:
@@
Let $N\geq 100$ be a positive integer, and $t\leq N$. Then we have the following results:
(i) For all $0<\epsilon<1$, if $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$, then
@@
(ii) There exist $f(N)\sim\frac{\sqrt{24N}}{\sqrt{\frac{6}{\pi}-1}}$, such that if $\frac{2\pi\sqrt{2N}}{\sqrt{\log N}}<t<f(N)$, then
@@
$Q_{t}(N)=\frac{2\sqrt{\pi}\exp\left(\frac{t-1}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}.$
(iii) For $t\geq f(N)$, where $f(N)$ is defined in part (ii), we have
$Z_{t}(N)\geq\frac{p(N)^{2}}{\exp\left(1.00873t\exp\left(-\frac{t(t-1)}{2\left(N+\frac{t^{2}-1}{24}\right)}\right)+\frac{2\pi}{\sqrt{6}}\frac{t}{\sqrt{N-t}+\sqrt{N}}\right)}\left(1+O\left(\frac{t}{N}\right)\right).$
@@ 2. Acknowledgments
We extend our heartfelt thanks to Prof. Ben Kane for his insightful and valuable suggestions. J. Barman is deeply thankful to the University Grants Commission (UGC), India, for their invaluable support through the Fellowship Programme. K. Mahatab is supported by the DST INSPIRE Faculty Award Program (grant no. DST/INSPIRE/04/2019/002586) and ARG-MATRICS program (grant no. ANRF/ARGM/2025/002540/MTR).
@@ 3. Preliminaries
If we consider the partitions $\mu$ having the largest part $t$ and the $t$-core partitions $\lambda$, then by the above lemma $\chi_{\lambda}(\mu)=0.$ This set of pairs $(\lambda,\mu)$ gives $c_{t}(N)p_{t}(N-t)$ zeros. We also notice that the set of partitions $\mu$ of $N$ having the largest part $t_{1}$ are disjoint from the set of partitions of $N$ having the largest part $t_{2}$. So the number of zeros in the character table is at least
@@
Next, we simplify Tyler’s formula for $c_{t}(N)$, which requires the following notations.
@@
where $z=x+iy$, $y>0$. In [14], Tyler defines the following functions:
$\mu_{k}(z)=-\frac{z^{k+1}}{2\pi i}\left(\frac{d}{dz}\right)^{k}\log\eta(z),\quad\text{for $k\geq 1$}$
$\text{and}\quad f_{t}(z)=\frac{\eta(tz)^{t}}{\eta(z)}.$
@@
For each $y\geq\frac{\sqrt{3}}{2}$, there exist a $v$ with $1<v<1.00873$ such that
$\eta(iy)=\exp\left(-\frac{\pi y}{12}-ve^{-2\pi y}\right).$
@@
(3.1) $\exp(-2\pi y)<\sum_{n=1}^{\infty}\frac{\sigma(n)}{n}\exp(-2\pi ny)<\sum_{n=1}^{\infty}n\exp(-2\pi ny)=\frac{\exp(-2\pi y)}{\left(1-\exp(-2\pi y)\right)^{2}}.$
Now we obtain an upper bound of the above choosing $y=\frac{\sqrt{3}}{2}$ and write
@@
For each $0<y<1$, there exist a $v$ with $1<v<1.00873$ such that
$\eta(iy)=y^{-\frac{1}{2}}\exp\left(-\frac{\pi}{12y}-ve^{-\frac{2\pi}{y}}\right).$
@@
$\eta\left(-\frac{1}{z}\right)=\sqrt{-iz}\,\eta(z),$
@@
Using Lemma 3.3, we obtain our required result. ∎
Lemma 3.5 ([14, Lemma 4.2]).
Let $\mu_{2}$ and $0<y\leq\frac{1}{10}$ be defined as before. Then for any positive integer $t>1$, the following inequalities hold:
$(i)$ For $ty<1$, we have
$\frac{2\sqrt{\pi}}{\sqrt{y(t-1)}}<\frac{1}{\sqrt{\mu_{2}(iy)-\mu_{2}(ity)}}<\frac{2\sqrt{2\pi}}{\sqrt{y(t-1)}}.$
$(ii)$ For $ty\geq 1$, we have
@@
Next, we recall the main theorem for $c_{t}(N)$ from Tyler’s paper [14].
Theorem 3.6 ([14, Theorem 1.4]).
Assume that $N\geq 0$ and $t\geq 6.$
(i) The equation
(3.2) $\frac{\mu_{1}(ity)-\mu_{1}(iy)}{y^{2}}=N+\frac{t^{2}-1}{24}$
admits a unique solution $y>0$, which satisfies
(3.3) $\frac{t-1}{4\pi\left(N+\frac{t^{2}-1}{24}\right)}<y<\frac{1}{\frac{3}{\pi}+\sqrt{24N-1+\frac{9}{\pi^{2}}}}.$
(ii) With this value of $y$, we have
(3.4) $c_{t}(N)=\frac{y^{\frac{3}{2}}\exp\left(2\pi y\left(N+\frac{t^{2}-1}{24}\right)\right)(\eta(ity))^{t}}{\sqrt{\mu_{2}(iy)-\mu_{2}(ity)}\,\eta(iy)}\left(1+O\left(\frac{1}{\text{min}(t,\sqrt{N})}\right)\right).$
Now, we simplify Tyler’s bound [14] for $c_{t}(N)$ in different ranges of $t$.
@@
Let $N\geq 100$ be a positive integer, and $t\leq N$.
(i) For all $0<\epsilon<1$, if $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$, then
@@
(ii) There exists $f(N)\sim\frac{\sqrt{24N}}{\sqrt{\frac{6}{\pi}-1}}$ such that, if $\frac{2\pi\sqrt{2N}}{\sqrt{\log N}}<t<f(N)$, then
$c_{t}(N)\geq\frac{2\sqrt{\pi}\exp\left(\frac{t-1}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}\left(1+O\left(t^{-1}\right)\right).$
(iii) For $t\geq f(N)$, where $f(N)$ is defined in part (ii), we have
$c_{t}(N)\geq p(N)\exp\left(-1.00873t\exp\left(-\frac{t(t-1)}{2\left(N+\frac{t^{2}-1}{24}\right)}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
(iv) For $\frac{\sqrt{6}}{2\pi}\sqrt{N}\log N<t$ and $N\geq 300000$, we have
$c_{t}(N)\geq p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right).$
Remark 1.
Part $(i)$ of the above proposition gives an asymptotic bound for $c_{t}(N)$. In parts $(ii)$ and $(iii)$, we separate different ranges of $t$ according to whether $ty<1$ or $ty\geq 1$.
Remark 2.
Part (iv) of Proposition 3.7 plays a central role in the proof of Theorem 1.2 as it provides a sharper bound than part (iii) for large NN. Here, we assume $N>300000>e^{4\pi}$ to ensure that $ty\geq\frac{\sqrt{N}}{2\pi}\sqrt{N}\log N\frac{1}{\sqrt{24N}}\geq 1$.
@@
$(i)$ From equation (1.7) of [14], when $6\leq t\leq\frac{2\pi\sqrt{2N}}{\sqrt{(1+\epsilon)\log N}}$ and $0<\epsilon<1$, we have
@@
Substituting the above in (3.5), we obtain
@@
$(ii)$ From (3.3) and the specified range of $t$, we observe that $ty<1$. By Lemma 3.4, we obtain
(3.6) $\eta(ity)\geq(ty)^{-\frac{1}{2}}\exp\left(-\frac{\pi}{12ty}-1.00873e^{-2\pi}\right).$
To obtain the following bound for $c_{t}(N)$, we consider the formula (3.4), and put in the lower bound for $\eta(ity)$ from (3.6), upper bound for $\eta(iy)$ from Lemma 3.4, and upper bound of $\mu_{2}(iy)-\mu_{2}(ity)$ from Lemma 3.5
@@
In the above bound, we make $y$ explicit in terms of $t$ and $N$ by minimizing the right-hand side at $y=\frac{t-1}{4\pi\left(N+\frac{t^{2}-1}{24}\right)}$, lower end point of the interval of $y$ in (3.3), and obtain
$c_{t}(N)\geq\frac{2\sqrt{\pi}\exp\left(\frac{t-1}{2}-1.00873te^{-2\pi}\right)\left(\frac{\pi}{6}(24N+t^{2}-1)\right)^{\frac{t-3}{2}}}{t^{t-1}}\left(1+O\left(t^{-1}\right)\right).$
$(iii)$ Again by (3.3), we see that $ty\geq 1$ for the specified range of $t$. From Lemma 3.3, we can write
(3.7) $\eta(ity)\geq\exp\left(-\frac{\pi ty}{12}-1.00873e^{-2\pi ty}\right).$
Similar to the proof of $(ii)$, we consider the asymptotic formula of $c_{t}(N)$ from Theorem 3.6 $(ii)$ and use the bounds of $\eta(ity)$, $\eta(iy)$, and $\mu_{2}(iy)-\mu_{2}(ity)$ from equation (3.7), Lemma 3.4 and Lemma 3.5 respectively. This gives
$c_{t}(N)\geq\sqrt{12}y^{2}\exp\left(2\pi Ny+\frac{\pi}{12y}-\frac{\pi y}{12}-1.00873t\exp(-2\pi yt)+e^{-\frac{2\pi}{y}}\right)$
$\hskip 256.0748pt\times\left(1+O\left(N^{-\frac{1}{2}}\right)\right)$
$\geq\sqrt{12}y^{2}\exp\left(\left(\sqrt{2\pi Ny}-\sqrt{\frac{\pi}{12y}}\right)^{2}+\frac{2\pi\sqrt{N}}{\sqrt{6}}-1.00873t\exp(-2\pi yt)+e^{-\frac{2\pi}{y}}\right)$
$\hskip 256.0748pt\times\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
To minimize the above expression, we set the squared term in the exponent equal to zero. From (3.3), we have
$\exp\left(-1.00873t\exp(-2\pi yt)\right)\geq\exp\left(-1.00873t\exp\left(-\frac{t(t-1)}{2\left(N+\frac{t^{2}-1}{24}\right)}\right)\right).$
Also, note
$\exp\left(e^{-\frac{2\pi}{y}}\right)=1+O\left(N^{-\frac{1}{2}}\right).$
Using these simplifications and the asymptotic result for the partition function (1.1), we have
$c_{t}(N)\geq p(N)\exp\left(-1.00873t\exp\left(-\frac{t(t-1)}{2\left(N+\frac{t^{2}-1}{24}\right)}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
$(iv)$ From Lemma 3.3, we have
(3.8) $\eta(ity)=\exp\left(-\frac{\pi ty}{12}-ve^{-2\pi ty}\right),$
where $1<v<\alpha=\frac{1}{(1-\exp(-2\pi yt))^{2}}$. The value of $\alpha$ comes from equation (3.1).
Using the explicit expression of $\mu_{1}$ for large***We refer to the imaginary part as large when it lies in $1,\infty)$, and as small when it lies in $(0,1)$. imaginary parts (see [14, (4.14) and (4.16)]), we deduce
$\mu_{1}(ity)=-\sum_{n=1}^{\infty}t^{2}y^{2}\sigma(n)\exp(-2\pi nty)+\frac{t^{2}y^{2}}{24},$
and for small imaginary parts
$\mu_{1}(iy)=\sum_{n=1}^{\infty}\sigma(n)\exp\left(-\frac{2\pi n}{y}\right)-\frac{1}{24}+\frac{y}{4\pi}.$
Substituting the above expressions for $\mu_{1}$ in (3.2), we obtain
(3.9) $y^{2}\left(N-\frac{1}{24}\right)+\frac{y}{4\pi}-\left(\frac{1}{24}-K_{1}-K_{2}\right)=0,$
where
$K_{1}=\sum_{n=1}^{\infty}t^{2}y^{2}\sigma(n)\exp(-2\pi nty)\,\,\text{and }K_{2}=\sum_{n=1}^{\infty}\sigma(n)\exp\left(-\frac{2\pi n}{y}\right).$
We may consider (3.9) as a quadratic equation and solve
$y=-\frac{1}{8\pi\left(N-\frac{1}{24}\right)}+\frac{1}{\sqrt{24\left(N-\frac{1}{24}\right)}}\sqrt{1+\frac{3}{8\pi^{2}\left(N-\frac{1}{24}\right)}-24(K_{1}+K_{2})}.$
Substituting the expression for $y$ into $K_{1}$ Note that $K_{1}=\sum_{n=1}^{\infty}t^{2}y^{2}\sigma(n)\exp(-2\pi nty)\leq\frac{(\log N)^{2}}{16\pi^{2}\sqrt{N}}+O\left(\frac{(\log N)^{2}}{N}\right)$ for the specified range of $t$. This estimate is used in the derivation of (3.10). and $K_{2}$ We also note that $K_{2}=\sum_{n=1}^{\infty}\sigma(n)\exp\left(-\frac{2\pi n}{y}\right)=O(N^{-2}).$ in the given range of $t$, and simplifying further, we obtain
(3.10) $y=\frac{1}{\sqrt{24N}}+\frac{C_{1}(N,t)}{N}+\frac{C_{2}}{N}+O\left(\frac{(\log N)^{2}}{N^{\frac{3}{2}}}\right)$
is the solution of (3.9), where $0<|C_{1}(N,t)|\leq C_{3}(\log N)^{2}$ and, $C_{2}$ and $C_{3}$ are constants independent of $N$. We may also compute
(3.11) $y^{-1}=\sqrt{24N}\left(1-\frac{C_{1}(N,t)\sqrt{24}}{\sqrt{N}}-\frac{C_{2}\sqrt{24}}{\sqrt{N}}+O\left(\frac{(\log N)^{2}}{N}\right)\right).$
We next use the asymptotic formula for $c_{t}(N)$ given in Theorem 3.6 $(ii)$ and combine it with the bounds for $\eta(ity)$, $\eta(iy)$, and $\mu_{2}(iy)-\mu_{2}(ity)$ from equation (3.8), Lemma 3.4, and Lemma 3.5, respectively. This yields
@@
(3.12) $\hskip 204.85974pt\times\left(1+O\left(N^{-\frac{1}{2}}\right)\right).$
Now we simplify the terms inside the exponential in the following steps. Note from (3.10) and (3.11)
(3.13) $2\pi Ny+\frac{\pi}{12y}=\frac{2\pi}{\sqrt{6}}\sqrt{N}+O\left(\frac{(\log N)^{2}}{N^{\frac{1}{2}}}\right).$
We also have
(3.14) $-tv\exp(-2\pi yt)$ $=-tv\exp(-2\pi y_{0}t)+tv\exp(-2\pi y_{0}t)\left(1-\exp(-2\pi tE)\right)$
where
$y_{0}=\frac{1}{\sqrt{24N}}\,\,\text{and }E=\frac{C_{1}(N,t)}{N}+\frac{C_{2}}{N}+O\left(\frac{(\log N)^{2}}{N^{\frac{3}{2}}}\right).$
We take $\exp$ of (3.14) and substitute the value of $y$ from (3.10) to get
$\exp\left(-tv\exp(-2\pi yt)\right)$
$=\exp\left(-tv\exp\left(-2\pi y_{0}t\right)\right)\exp\big(tv\exp(-2\pi y_{0}t)\big(1-\exp(-2\pi tE)\big)\big)$
(3.15) $=\exp\left(-tv\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right).$
Here we used the fact that in the range $\frac{\sqrt{6}}{2\pi}\sqrt{N}\log N<t\leq N$, the error attains its maximum at $t=\frac{\sqrt{6}}{2\pi}\sqrt{N}\log N$, which gives
$\exp\big(tv\exp(-2\pi y_{0}t)\big(1-\exp(-2\pi tE)\big)\big)=1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right).$
Now, using simplifications of (3.10), (3.13) and (3) in (3), and using asymptotic formula for $p(N)$ from (1.1), we obtain
$c_{t}(N)$ $\geq\frac{\exp\left(\frac{2\pi}{\sqrt{6}}\sqrt{N}\right)}{4\sqrt{3}N}\exp\left(-tv\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right)$
$=p(N)\exp\left(-tv\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right).$
From (3.8), $1<v<\alpha=\frac{1}{(1-\exp(-2\pi yt))^{2}}$. One can easily check that for the given range of $t$, $\alpha=1+O\left(N^{-\frac{1}{2}}\right)$. Consequently,
$c_{t}(N)\geq p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right).$
In the next lemma, we present a sketch of the proof of the Erdös and Lehner’s result for $p_{t}(N)$ [3], with an explicit error term, which is an addition to the original paper.
@@
where $C=2\pi/\sqrt{6}$ and $x\ll N^{\frac{1}{4}-\gamma},\ \gamma>0$.